1. You own a portfolio invested 17.94% in Stock A, 10.6% in Stock B, 13.95% in Stock C, and the remainder in Stock D. The beta of these four stocks are 1.07, 0.93, 0.21, and 1.45. What is the portfolio beta?
Answer #1

Weight of stock D = 100 - (17.94 + 10.6 +13.95) = 100 - 42.49 = 57.51 %

Portfolio beta, ßP = WA x ßA+ WB x ßB + WC x ßC + WD x ßD

WA, WB, WC, WD = Weight of stock A, B, C, D = 17.94 %, 10.6 %, 13.95 %, 57.51 % respectively

ßA, ßB, ßC, ßD= Beta of stock A, B, C, D = 1.07, 0.93, 0.21, 1.45 respectively

ßP = 0.1794 x 1.07+ 0.106 x 0.93 + 0.1395 x 0.21 + 0.5751 x 1.45

    = 0.191958 + 0.09858 + 0.029295 + 0.833895 = 1.153728

Beta of the portfolio is 1.153728